You want to be able to withdraw $2500 per month during your retirement. You want to be able to do this for 25 years, and your account will earn 8% interest compounded monthly.

How much do you need to have in your account at the beginning of retirement?

How much total money will you pull out of the account during the 25 years?

How much of that money is interest?
Answer #1

1.
=2500/(8%/12)*(1-1/(1+8%/12)^(12*25))=323911.3065

2.
=2500*12*25=750000

3.
=750000-323911.3065=426088.6935